OpenStudy (anonymous):

An unknown compound contains only carbon, hydrogen, and oxygen (). Combustion of 7.50 of this compound produced 11.0 of carbon dioxide and 4.50 of water. 1) How many moles of carbon, C, were in the original sample? An unknown compound contains only carbon, hydrogen, and oxygen (). Combustion of 7.50 of this compound produced 11.0 of carbon dioxide and 4.50 of water. 1) How many moles of carbon, C, were in the original sample? @Chemistry

OpenStudy (anonymous):

n(C)=m/M m(CO2)=11.0 M(CO2)=44.01 n=11/44.01 =0.249943194moles = moles of Carbon If you wanted to find the empirical formula though, n(H)=4.5/18.016 =0.249777975x2= 0.49955595moles of H m(H)= 0.499556x1.008=0.50355 m(C)= 0.2499x12.01= 3.0018 7.50-0.50355-3.0018 = 3.99g = 0.24966moles of O C =0.2499/0.2497 =1 H = 0.4996/0.2497 =2 0 = 0.2497/0.2497 =1 therefore the empirical formula=CH2O